While walking home late at night, a woman was knocked over and raped. He didn’t leave...

  1. While walking home late at night, a woman was knocked over and raped. He didn’t leave any hairs at the crime scene but left eyeglasses with unusual frames. An optician helped police locate him. The man also left a half-eaten lollipop at the scene. The DNA taken from the suspect matched that found on the lollipop at four STRs on different chromosomes. Allele frequencies from the man’s ethnic group are listed besides the prolife pattern show here:
    • STR 1         Homozygote     Allele frequency = 0.2
    • STR 2         Allele 1 = 0.3    Allele 2 = 0.7
    • STR3          Homozygote     Allele frequency = 0.1
    • STR4          Allele 1 = 0.4     Allele 2 = 0.2
  1. How would the electropherogram for STRs 1 and 3 look different from those for 2 and 4?
  2. What is the probability that the suspect’s DNA matches that of the lollipop rapist by chance? Show the calculations
  3. The man’s population group is highly inbred. How does this affect the accuracy or reliability of the DNA profile? (P.S. the suspect was so frightened by the DNA analysis that he confessed!)
Answer #1

a) Electropherogram is the chart produced when the electrophoresis is used in an analytical technique.

Refer the below fig. (Source: Creator:Butler, John M.Information extracted from IPTC Photo Metadata)

Observed profile A B U M L ULUL 4 alleles All heterozygotes and non-overlapping alleles EN ALL A 3 alleles Heterozygote + het

STR 1 and 3 will give a single band while 2 and 4 have two alleles so two bands are expected.

b)

STR 1 and 3 is matched between two individuals.Thus the probability matching is 1.

The probability of STR2 and STR4 to be allele 1 is 0.3 or allele2 is 0.7. Therefore there is 1/2 probability for the occurrence of each allele.

the probability of both suspect and man having same allele is 1/2 * 1/2 = 1/4 in both cases of STR 2 & STR 4

Thus the probability of matching is 1*1/4*1*1/4 = 1/16 (Using "and rule" of probability)

c) Inbred population is tend to have same allele for a gene. It leads to generate identical population. Thus it may affect the accuracy of the DNA profile.